# Adaptive Rank, Reduced Forgetting: Knowledge Retention in Continual Learning Vision-Language Models with Dynamic Rank-Selective LoRA

### v6

@@ 3.2 Analyses on LoRA Location and Rank
To enable efficient and natural continual updating of PTMs, we employ LoRA and analyze its learning and forgetting capability for optimizing its usage. We examine how LoRA’s rank and placement affect downstream learning and forgetting through two types of analyses (Pearl, 2022; Meng et al., 2022), providing a comprehensive view of their overall impact. (1) We first apply LoRA components with varying ranks at different locations in the pre-trained CLIP, for measurement of importance or contribution (Fig. 4); (2) We apply LoRA across all PTM weights and systematically prune them for evaluation, for testing the standalone capability (Fig. 4). We train the model on downstream datasets and assess its performance on both (1) the new task, i.e., learning new knowledge, and (2) an additional reference dataset to evaluate the retention and forgetting of existing capabilities. We achieve three main results to guide methodology design. Without loss of generality, the learning–forgetting analyses can be extended to each learning step in straightforward, assumption-free CL.
@@
Applying LoRA on all weights. Unlike prior works (Hu et al., 2021; Zhang et al., 2024a; Meng et al., 2024) that only apply LoRA to limited components by default, we first examine the impact of different LoRA placements. As shown in Fig. 4 and Fig. 4, applying LoRA to only specific modules alone may limit learning capacity in different ways: (1) LoRA placement strongly impacts downstream performance and retention of pre-trained capabilities; (2) Updating only the vision encoder causes more forgetting; (3) Updating only attention modules mitigates forgetting but limits learning on new tasks. These observations reflect the distinct roles of PTM components. Takeaway 1: Instead of manual or fixed LoRA placement at specific parameters of PTM like previous methods (Hu et al., 2021; Liang & Li, 2024; Yu et al., 2024; Zhang et al., 2024a; Meng et al., 2024), we apply LoRA to all weights and optimize for an adaptive configuration.
@@
The effects of LoRA ranks on balancing learning and forgetting. We identified that the rank number of LoRA plays a critical role in balancing learning and forgetting. As shown in Fig. 4, we can observe that a relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while a lower-rank LoRA mitigates forgetting but limits adaptation. This suggests that lower-rank LoRA interferes less with existing knowledge (i.e., reduces forgetting) while learning less new knowledge. Fortunately, we identify a balance (top-right in Fig. 4), where a moderately small rank reduces forgetting while enabling enough adaptability.
@@
During training, low-rank parameter updates tend to enable the updated model to be closer to the original model, with less interference and forgetting. In the extreme case, a rank-zero LoRA learns nothing and introduces no forgetting. Thus, minimizing LoRA rank for each new task in CL can reduce forgetting without extra assumptions. On the other hand, the task-learning objective emphasizes new knowledge acquisition, enabling a balance. Takeaway 2: A plasticity–stability balance exists and associates with LoRA rank (high vs. low), which can be adaptively achieved by jointly optimizing the task objective and minimizing LoRA ranks.
@@
The balance varies by location and task. Fig. 4 shows that the optimal rank for balance depends on both parameter placement and the learning task, consistent with the pruning-based sufficiency analysis in Fig. 4. For example, removing updates to the Value projection in the Attention module of the vision encoder had little impact on adaptation but substantially restored pre-trained performance. Interestingly, pruning the MLP layer slightly reduced task performance while enhancing pre-trained capabilities. Takeaway 3: The learning–forgetting balance point tied to rank varies across modules and tasks, necessitating adaptive optimization.
@@ 3.3 Proposed CoDyRA with Sparsity-Induced Rank Selection and Minimization
We introduce adaptive rank-selective updates for all pre-trained weight matrices in the Attention and MLP modules of both vision and text encoders (Fig. 2). The sparsity-promoting regularization enables adaptive rank selection and minimization for each LoRA adapter on each task. For each task in CL, beyond adaptive rank minimization, CoDyRA requires no additional priors or assumptions and does not store or maintain information (e.g., data, parameters, or statistics) from previous steps, unlike prior methods (Zheng et al., 2023; Yu et al., 2024; Xu et al., 2024).
@@
Regularized Low-rank updates with rank importance. To update weight matrix $\mathbf{W}_{0}^{t,m}\in\mathbb{R}^{d\times k}$, LoRA (Hu et al., 2021) introduces two low-rank matrices, $\mathbf{B}^{t,m}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}^{t,m}\in\mathbb{R}^{r\times k}$, such that $\Delta\mathbf{W}^{t,m}=\mathbf{B}^{t,m}\mathbf{A}^{t,m}$, where $r\ll\min(d,k)$. The updated weight matrix is then defined as:
@@
Adaptive rank-selective updates with sparse-prompting regularization. The formulation in Eq. (2) enables the model to dynamically optimize the importance of each rank and minimize active ranks through gradient descent. [^1] To align the training process with the pre-trained model, we use CLIP’s contrastive loss as the optimization objective during pre-training. To only retain the essential ranks for learning new tasks (thus reducing forgetting), we can achieve optimizing $R(\Delta\mathbf{W}^{t,m})$ through an $\ell_{1}$ norm-based sparsity-promoting regularization on $\mathbf{w}^{t,m}$, i.e., $\|\mathbf{w}^{t,m}\|_{1}$. The optimization objective for parameters $\{\mathbf{w}^{t,m},\mathbf{B}^{t,m},\mathbf{A}^{t,m}\}_{m=1}^{M}$ learned on task $t$ is:
$\mathcal{L}_{\text{train}}^{t}:={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{L}_{\text{sup}}^{t}}}+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\lambda\sum\nolimits_{m=1}^{M}\|\mathbf{w}^{t,m}\|_{1}}},$ (3)
where $\mathcal{L}_{\text{sup}}^{t}$ is the supervised training loss for current task, $M$ is the number of weight matrices inserted $r$-rank update, and $\mathbf{w}^{t,m}$ represents the importance weights of each rank added to the weight matrix $\mathbf{W}^{t,m}_{0}$. The $\ell_{1}$ regularization strength is controlled by $\lambda$.
To handle the non-differentiable $\ell_{1}$ regularization applied to the importance weights, we adopt the proximal gradient method (Beck & Teboulle, 2009). The $i$-th element of $\hat{\mathbf{w}}^{t,m}$ for rank $i$ is updated via soft-thresholding:
$\mathbf{w}_{i}^{t,m}:=\mathbbm{1}(\lvert\hat{\mathbf{w}}^{t,m}_{i}\rvert>\kappa)\cdot(\hat{\mathbf{w}}^{t,m}_{i}+\mathrm{sign}(\hat{\mathbf{w}}^{t,m}_{i})\cdot\kappa),$ (4)
@@
This approach preserves only the ranks with significant importance for learning the current task and prunes those with low relevance, adapting across weights and tasks. The importance weights on preserved ranks softly highlight their relative significance. Early in training, dense updates are applied without Eq. (4), allowing all ranks to capture task-relevant information before sparsity is enforced.
Continually learn on new tasks. CoDyRA views CL as a sequence of tasks with LoRA, without storing past data Zheng et al. (2023), task information Yu et al. (2024); Xu et al. (2024), or relying on assumptions (Liang & Li, 2024). For each new task, new CoDyRA modules are initialized and optimized; low-importance ranks are pruned, and non-zero ranks merged into the weights. Sparsity-promoting regularization controls task capacity, keeping the model close to its previous state and reducing forgetting. This enables efficient continual updating with only sparsity-based regularization.
Table 1: Comparisons on X-TAIL for each domain. The best and the second best results are highlighted in red and blue, respectively. Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection (Yu et al., 2024; Xu et al., 2024).
Method
@@
   Zero-shot (Radford et al., 2021) 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
   LwF (Li & Hoiem, 2017) 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
   WiSE-FT (Wortsman et al., 2022) 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
   iCaRL (Rebuffi et al., 2017b) 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
   ZSCL (Zheng et al., 2023) 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
[…]
   CoDyRA 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
[…]
Average
   LwF (Li & Hoiem, 2017) 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
   WiSE-FT (Wortsman et al., 2022) 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
   iCaRL (Rebuffi et al., 2017b) 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
   ZSCL (Zheng et al., 2023) 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
[…]
   CoDyRA 41.4 81.0 58.7 77.8 83.4 84.6 64.5 90.4 67.2 64.4 71.3
[…]
Last
   LwF (Li & Hoiem, 2017) 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   WiSE-FT (Wortsman et al., 2022) 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
   iCaRL (Rebuffi et al., 2017b) 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   ZSCL (Zheng et al., 2023) 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
[…]
   CoDyRA 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2
[…]
4 Experiments
4.1 Experimental Setting
Datasets. We evaluate our method on two widely used continual learning settings for vision-language models: MTIL (Zheng et al., 2023; Yu et al., 2024) and X-TAIL (Xu et al., 2024). A wide range of fine-grained datasets is used, where each dataset is treated as an individual task to be incrementally learned. We leave detailed experimental settings in Appendix B.
Evaluation Metrics. We use the same metrics as prior work (Zheng et al., 2023; Yu et al., 2024; Xu et al., 2024): “Transfer” for zero-shot performance on unseen data, “Last” for retention of earlier tasks, and “Average” for mean accuracy across all datasets of all learning tasks.
Implementation Details. We follow the setups in (Zheng et al., 2023; Yu et al., 2024; Xu et al., 2024) using CLIP with a ViT-B/16 backbone (Radford et al., 2021). By default, CoDyRA is applied to all pre-trained weight matrices in the vision and text encoders with an initial rank of 16. Each task is trained for 500 iterations with AdamW (Loshchilov & Hutter, 2017), using dense training ratios of 0.5 for X-TAIL and 0.7 for MTIL. $\kappa_{\text{max}}$ is set as 0.005.
4.2 Experimental Results
Cross-domain task-agnostic incremental learning. We report X-TAIL results in Table 1. Unlike prior methods (Yu et al., 2024; Xu et al., 2024) that rely on zero-shot predictions to preserve pre-trained capabilities, our approach continually updates the model with minimal forgetting. As those methods are bounded by zero-shot performance, ours surpasses this limit, achieving higher transfer accuracy. While continual updates may slightly affect some unseen predictions, the overall effect is positive: the model adapts to new tasks while modestly improving generalization.
[…]
Table 2: Average transfer accuracy on unseen datasets not encountered during CL.
Method Average Transfer Accuracy on Unseen Data
CIFAR100 Places365 ImageNet-1k Average
Zero-Shot (Radford et al., 2021) 68.24 33.77 66.72 56.24
MoE-Adapter (Yu et al., 2024) 68.24 33.77 66.72 56.24
RAIL (Xu et al., 2024) 68.24 33.77 66.72 56.24
CoDyRA 68.95 36.52 68.52 58.00
Table 3: Comparisons on 5-shot MTIL setting.
Method
Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
CLIP
   Zero-shot (Radford et al., 2021) 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.4 89.0 64.7 65.2 65.3
Transfer
   Zero-shot (Radford et al., 2021) 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
   LwF (Li & Hoiem, 2017) 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
   LwF-VR (Ding et al., 2022) 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
   WiSE-FT (Wortsman et al., 2022) 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
   ZSCL (Zheng et al., 2023) 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
@@ 4.2 Experimental Results
   LwF (Li & Hoiem, 2017) 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
   LwF-VR (Ding et al., 2022) 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
   WiSE-FT (Wortsman et al., 2022) 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
   ZSCL (Zheng et al., 2023) 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
@@
   LwF (Li & Hoiem, 2017) 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
   LwF-VR (Ding et al., 2022) 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
   WiSE-FT (Wortsman et al., 2022) 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
   ZSCL (Zheng et al., 2023) 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
@@
Multi-domain task-incremental learning. We further evaluate our method in the few-shot MTIL setting (Table 3), following (Yu et al., 2024; Xu et al., 2024). As with X-TAIL, our approach outperforms prior methods across all metrics without extra techniques and even surpasses their zero-shot upper bound in average transfer accuracy. Incorporating techniques from (Yu et al., 2024; Xu et al., 2024) further boosts performance.

### v7

@@ 3.2 Analyses on LoRA Location and Rank
To enable efficient and natural continual updating of PTMs, we employ LoRA and analyze its learning and forgetting capability for optimizing its usage. We examine how LoRA’s rank and placement affect downstream learning and forgetting through two types of analyses [59, 54], providing a comprehensive view of their overall impact. (1) We first apply LoRA components with varying ranks at different locations in the pre-trained CLIP, for measurement of importance or contribution (Fig. 3); (2) We apply LoRA across all PTM weights and systematically prune them for evaluation, for testing the standalone capability (Fig. 3). We train the model on downstream datasets and assess its performance on both (1) the new task, i.e., learning new knowledge, and (2) an additional reference dataset to evaluate the retention and forgetting of existing capabilities. Without loss of generality, the learning–forgetting analyses can be extended to each learning step in straightforward, assumption-free CL.
@@
Applying LoRA on all weights. Unlike prior works [27, 95, 53] that only apply LoRA to limited components by default, we first examine the impact of different LoRA placements. As shown in Fig. 3 and Fig. 3, applying LoRA to only specific modules alone may limit learning capacity in different ways: (1) LoRA placement strongly impacts downstream performance and retention of pre-trained capabilities; (2) Updating only the vision encoder causes more forgetting; (3) Updating only attention modules mitigates forgetting but limits learning on new tasks. These observations reflect the distinct roles of PTM components. Takeaway 1: Instead of manual or fixed LoRA placement [27, 95, 53], we apply LoRA to all weights and optimize for an adaptive configuration.
@@
The effects of LoRA ranks on balancing learning and forgetting. We identified the rank number of LoRA plays a critical role for balancing learning and forgetting. As shown in Fig. 3, we can observe that a relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while a lower-rank LoRA mitigates forgetting but limits adaptation. This suggests that lower-rank LoRA interferes less with existing knowledge (i.e., reduces forgetting) while learning less new knowledge. Fortunately, we identify a balance (top-right in Fig. 3), where a moderately small rank reduces forgetting while enabling enough adaptability.
@@
During training, low-rank updates tend to enable the updated model to be closer to the original model, with less interference and forgetting. In the extreme case, a rank-zero LoRA learns nothing and introduces no forgetting. Thus, minimizing LoRA rank for each new task in CL can reduce forgetting without extra assumptions. On the other hand, the task-learning objective emphasizes new knowledge acquisition, enabling a balance. Takeaway 2: A plasticitystability balance exists and associates with LoRA rank (high vs. low), which can be adaptively achieved by jointly optimizing the task objective and minimizing LoRA ranks.
@@
The balance varies by location and task. Fig. 3 shows that the optimal rank for balance depends on both parameter placement and the learning task, consistent with the pruning-based sufficiency analysis in Fig. 3. For example, removing updates to the Value projection in the Attention module of the vision encoder had little impact on adaptation but substantially restored pre-trained performance. Interestingly, pruning the MLP layer slightly reduced task performance while enhancing pre-trained capabilities. Takeaway 3: The learningforgetting balance point tied to rank varies across modules and tasks, necessitating adaptive optimization.
@@ 3.3 Proposed CoDyRA with Sparsity-Induced Rank Selection and Minimization
We introduce adaptive rank-selective updates for all pre-trained weight matrices in the Attention and MLP modules of both vision and text encoders (Fig. 4). The sparsity-promoting regularization enables adaptive rank selection and minimization for each LoRA adapter on each task. For each task in CL, beyond adaptive rank minimization, CoDyRA requires no additional priors or assumptions and does not store or maintain information (e.g., data, parameters, or statistics) from previous steps, unlike prior methods [99, 93, 90].
@@
Regularized low-rank updates with rank importance. To update weight matrix $\mathbf{W}_{0}^{t,m}\in\mathbb{R}^{d\times k}$, LoRA [27] introduces two low-rank matrices, $\mathbf{B}^{t,m}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}^{t,m}\in\mathbb{R}^{r\times k}$, such that $\Delta\mathbf{W}^{t,m}=\mathbf{B}^{t,m}\mathbf{A}^{t,m}$, where $r\ll\min(d,k)$. The updated weight matrix is defined as:
@@
Adaptive rank-selective updates with sparse-prompting regularization. The formulation in Eq. (2) enables the model to dynamically optimize the importance of each rank and minimize active ranks through gradient descent. To only retain the essential ranks for learning new tasks (thus reducing forgetting), we can achieve optimizing $R(\Delta\mathbf{W}^{t,m})$ through an $\ell_{1}$ norm-based sparsity-promoting regularization [71, 14] on $\mathbf{w}^{t,m}$, i.e., $\|\mathbf{w}^{t,m}\|_{1}$. The optimization objective for parameters $\{\mathbf{w}^{t,m},\mathbf{B}^{t,m},\mathbf{A}^{t,m}\}_{m=1}^{M}$ learned on task $t$ is:
$\mathcal{L}_{\text{train}}^{t}:={{\color[rgb]{0.55,0.14,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.55,0.14,1}\pgfsys@color@cmyk@stroke{0.45}{0.86}{0}{0}\pgfsys@color@cmyk@fill{0.45}{0.86}{0}{0}\mathcal{L}_{\text{sup}}^{t}}}+{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\lambda\sum\nolimits_{m=1}^{M}\|\mathbf{w}^{t,m}\|_{1}}},$ (3)
where $\mathcal{L}_{\text{sup}}^{t}$ is the supervised training loss for current task, $M$ is the number of weight matrices inserted $r$-rank update, and $\mathbf{w}^{t,m}$ represents the importance weights of each rank added to the weight matrix $\mathbf{W}^{t,m}_{0}$. The $\ell_{1}$ regularization strength is controlled by $\lambda$. Minimizing $\|\mathbf{w}^{t,m}\|_{1}$ encourages minimizing of $\|\Delta\mathbf{W}^{t,m}\|_{F}$ if given $\mathbf{A}^{t,m}$ and $\mathbf{B}^{t,m}$ (according to Eq. (2)). It softly reduces capacity in updates conflicting with critical information encoded in the learned parameters, without introducing additional assumptions. More discussions are in Appendix A.3.
To handle the non-differentiable $\ell_{1}$ regularization applied to the importance weights, we adopt the proximal gradient method [4]. The $i$-th element of $\hat{\mathbf{w}}^{t,m}$ for rank $i$ is updated via soft-thresholding:
$\mathbf{w}_{i}^{t,m}:=\mathbbm{1}(\lvert\hat{\mathbf{w}}^{t,m}_{i}\rvert>\kappa)\cdot(\hat{\mathbf{w}}^{t,m}_{i}{-}\mathrm{sign}(\hat{\mathbf{w}}^{t,m}_{i})\cdot\kappa),$ (4)
@@
This approach preserves only the ranks with significant importance for learning current task and prunes those with low relevance, adapting across weights and tasks. The importance weights on preserved ranks softly highlight their relative significance. Early in training, dense updates are applied without Eq. (4), allowing all ranks to capture task-relevant information before sparsity is enforced.
Table 1: Comparisons on 5-shot MTIL setting. The best and the second best results are highlighted in red and blue, respectively. Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection [93, 90].
    Method Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
@@
   Zero-shot [62] 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
   LwF [38] 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
   LwF-VR [15] 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
   WiSE-FT [85] 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
   ZSCL [99] 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
@@
   LwF [38] 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
   LwF-VR [15] 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
   WiSE-FT [85] 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
   ZSCL [99] 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
@@
   LwF [38] 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
   LwF-VR [15] 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
   WiSE-FT [85] 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
   ZSCL [99] 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
@@
Continual learning framework. For each task, we initialize and optimize CoDyRA modules, low-importance ranks are pruned via sparsity-promoting regularization, and remaining active ranks are merged into the backbone weights after training. This regularization naturally constrains update capacity, preserving prior knowledge while enabling efficient adaptation. Following standard conventions, we initialize LoRA matrix $\mathbf{A}^{t,m}$ randomly and $\mathbf{B}^{t,m}$ to zero, and our proposed importance $\mathbf{w}^{t,m}$ is randomly initialized.
4 Experiments
We evaluate our approach on continual learning tasks for both vision-language models and LLMs. For vision-language CLIP [62, 30] model, we utilize the MTIL [99] and X-TAIL [90] benchmarks. For LLMs, we assess performance on the TRACE benchmark [81] using LLaMA [72, 16] and Gemma [70]. Detailed experiment settings and additional results are provided in Appendix B.
4.1 Experimental Results
[…]
Experiments on language model continual learning. Table 2 evaluates CoDyRA on the TRACE benchmark [81] using LLaMA [16, 72] and Gemma [70]. Traditional regularization (EWC [33], OGD [18]), rehearsal (GEM [44]), and prompt-based methods (L2P [84], DualPrompt [83]) often underperform simple In-Context Learning (ICL [7]). Among parameter-efficient approaches, naive SeqLoRA suffers from significant forgetting. While O-LoRA [80], HiDeLoRA [76, 77], and TreeLoRA [60] improve stability, they impose rigid orthogonality, complex decomposition, or structural expansion. In contrast, CoDyRA achieves state-of-the-art performance without such constraints. By leveraging adaptive rank selection via sparsity-promoting regularization, it dynamically allocates capacity (minimizing ranks for stability while expanding for plasticity) to optimize the trade-off between learning and forgetting.
Table 2: Comparison with a broad range of CL methods on the TRACE benchmark. We report Overall Performance (OP (%) $\uparrow$) and Backward Transfer (BWT (%) $\downarrow$). Results are averaged over three runs with standard deviations. The best results are highlighted in bold.
FIX(ICL) SeqLoRA OGD GEM EWC L2P DualPrompt HiDeLoRA O-LoRA TreeLoRA CoDyRA
meta-llama / LLaMA-2-7B-Chat
OP 38.94 $\pm$ 0.3 34.3 $\pm$ 1.2 42.09 $\pm$ 1.6 40.08 $\pm$ 1.6 42.36 $\pm$ 1.2 36.23 $\pm$ 0.8 37.69 $\pm$ 1.2 41.60 $\pm$ 0.8 42.78 $\pm$ 0.8 43.52 $\pm$ 1.0 43.82 $\pm$ 0.9
BWT 18.5 $\pm$ 0.8 8.06 $\pm$ 1.2 6.77 $\pm$ 1.2 5.97 $\pm$ 0.8 8.25 $\pm$ 0.8 8.03 $\pm$ 0.8 7.12 $\pm$ 0.4 7.16 $\pm$ 0.4 3.46 $\pm$ 0.4 3.25 $\pm$ 0.5
google / Gemma-2B-it
OP 32.3 $\pm$ 0.2 31.89 $\pm$ 0.8 32.85 $\pm$ 1.4 26.48 $\pm$ 1.5 28.35 $\pm$ 1.6 31.14 $\pm$ 1.2 32.42 $\pm$ 1.0 33.25 $\pm$ 0.9 33.73 $\pm$ 0.8 33.41 $\pm$ 0.9 33.96 $\pm$ 1.0
BWT 15.28 $\pm$ 0.4 12.27 $\pm$ 0.9 18.25 $\pm$ 0.9 16.96 $\pm$ 1.2 15.77 $\pm$ 0.7 14.25 $\pm$ 0.5 13.66 $\pm$ 0.5 12.36 $\pm$ 0.4 8.50 $\pm$ 0.5 7.94 $\pm$ 0.8
meta-llama / LLaMA-3-1B-Instruct
OP 31.16 $\pm$ 0.4 29.73 $\pm$ 1.6 30.12 $\pm$ 2.0 32.19 $\pm$ 2.0 31.96 $\pm$ 1.6 29.38 $\pm$ 1.2 30.76 $\pm$ 1.2 33.73 $\pm$ 1.2 32.94 $\pm$ 0.8 36.14 $\pm$ 0.7 37.46 $\pm$ 0.8
BWT 17.03 $\pm$ 1.2 15.2 $\pm$ 1.6 10.74 $\pm$ 1.6 11.62 $\pm$ 1.2 13.57 $\pm$ 0.8 11.34 $\pm$ 0.8 12.36 $\pm$ 0.8 12.89 $\pm$ 1.2 7.36 $\pm$ 0.8 5.11 $\pm$ 0.8
Table 3: Pre-trained retention on unseen datasets.
Method Zero-shot Retention on Unseen Data
CIFAR100 Places365 ImageNet-1k Average
Zero-Shot [62] 68.24 33.77 66.72 56.24
MoE-Adapter [93] 68.24 33.77 66.72 56.24
RAIL [90] 68.24 33.77 66.72 56.24
CoDyRA 68.95 36.52 68.52 58.00